OpenStudy (anonymous):

Let f(x)=x^3-x-1 A. Find the equation of the tangent line to the graph of the function at x=0

OpenStudy (cwrw238):

find f'(x) in terms of x and put x = 0 this will give you the slope of the tangent line

OpenStudy (anonymous):

3x^2-1

OpenStudy (cwrw238):

so when x = 0 slope is 3(0) - 1 = -1

OpenStudy (anonymous):

that is the answer -1 I would write the answer as follow so when x=0 slope is 3(0)-1=-1

OpenStudy (cwrw238):

no -1 is the slope of the line use the general form of a straight line y-y1 = m(x-x1) m = slope = -1 and (x1,y1) is the coordinates of the point of contact of the tangent line we know the x1 = 0 so we need to find the value of y1 how would we do that?

OpenStudy (anonymous):

I am not sure really

OpenStudy (cwrw238):

just put x = 0 in x^3 -x-1

OpenStudy (anonymous):

6x-1

OpenStudy (cwrw238):

?? hoe did you get 6x??

OpenStudy (anonymous):

3x^2

OpenStudy (cwrw238):

0^3 - 0 -1 = ?

OpenStudy (anonymous):

4

OpenStudy (cwrw238):

how did you get 4? 0^3 = 0

OpenStudy (anonymous):

see I really don't know what I am doing

OpenStudy (anonymous):

y=0

OpenStudy (cwrw238):

y = 0 - 0 -1 = -1

OpenStudy (cwrw238):

now insert m = -1 , x1 = 0 and y1 = -1 into the equation to get the required equation

OpenStudy (cwrw238):

y - y1 = m(x - x1) y - (-1) = -1(x - 0)

OpenStudy (cwrw238):

now simplify

OpenStudy (anonymous):

this is what i get out of this y=-1=-1 i know this isn't right